OpenStudy (i_love_my_nieces):

Which ordered pair is not a solution of y = -4x -9

OpenStudy (i_love_my_nieces):

(-9, 27) (3, -21) (-6 15) (5, -28)

OpenStudy (shamim):

d

OpenStudy (shamim):

because y=-4x-9 -28=4*5-9 is not true

OpenStudy (shamim):

-28=-4*5-9 is not true

OpenStudy (shamim):

-28=-29